A population proportion is .40. A simple random sample of size 200 will be taken and the
sample proportion will be used to estimate the population proportion.
A. Compute the standard error of the proportion
B. What is the probability that the sample proportion will be within+.03 of the
population proportion?
C. What is the probability that the sample proportion will be within +.05 of the population
proportion?

Answer by Boreal(15235)

SE is sqrt(p*(1-p)/n)=sqrt(0.4*0.6/200)=0.0346
to be within 0.03, the z value would have to be +/-0.8670, which is probability 0.6140
to be within 0.05, the z value would have to be +/-1.445, which is probability 0.8515
z*SE=the margin of error.